If you earn a real rate of return of 2% and inflation is 4%, what is your nominal rate of return? (to 2 decimal places)
Answer #1

Nominal Rate of Return = ( 1+ Real rate)*(1+Inflation) -1 = (1+2%)*(1+4%) -1 = 6.08%

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well